Calculate the following binomial probability by either using one of the binomial probability tables, software, or a calculator using the formula below. Round your answer to 3 decimal places.

P(x | n, p) =

n!
(nx)! x!

· px · qnx    where    q = 1 − p

P(x = 11, n = 13, p = 0.70) =
Answer #1

Solution :

Given that x = 11 , n = 13 , p = 0.70

=> q = 1 - p = 0.30

=> For binomial distribution , P(x) = nCx*p^x*q^(n-x)

=> nCx = n!/((n-x)!*x!)

=> P(x = 11) = 13C11*0.70^11*0.30^2

= 0.1388

= 0.139 (rounded)